The 80th term of fibonacii sequence is?
1
Expert's answer
2021-09-28T12:58:17-0400

Use Binet formula


"F_n=\dfrac{\alpha^n-\beta^n}{\alpha-\beta}"

"\alpha=\dfrac{1+\sqrt{5}}{2}, \beta=\dfrac{1-\sqrt{5}}{2}"

"\alpha-\beta=\sqrt{5}"

"F_{80}=\dfrac{(1+\sqrt{5})^{80}-(1-\sqrt{5})^{80}}{2^{80}\sqrt{5}}"

"F_{80}= 14472334024676221"